OpenStudy (anonymous):

-6(5-2z) + 5z = 8z - 3(16-3z) Solve. Please show all your work?

OpenStudy (mertsj):

Why did you repost this?

OpenStudy (anonymous):

Because I ask someone else and they totally gave me a different answer to what you have given me

OpenStudy (mertsj):

What did they get?

OpenStudy (anonymous):

-6(5-2z) + 5z = 8z - 3(16-3z) -30 + 12z + 5z = 8z -48 +6z 17z= - 18 + 14z 3z = -18 z= -6

OpenStudy (mertsj):

There is just one problem with that solution: 3 times 3 is NOT 6

OpenStudy (anonymous):

So he is wrong?

OpenStudy (anonymous):

Can you please repost what you have given me?

OpenStudy (mertsj):

Do you not see? -3(16-3z)= -48+9z And he wrote that it is -48+6z

OpenStudy (anonymous):

Sorry about that.

OpenStudy (anonymous):

Can you please re post it again?

OpenStudy (mertsj):

−6(5−2z)+5z=8z−3(16−3z) −30+12z+5z=8z−48+9z 17z−30=17z−48 −30=−48

OpenStudy (anonymous):

I got -78 when I did -39=-48?

OpenStudy (mertsj):

there is NO SOLUTION Why don't you pick any number you want and plug it in in place of z and see what you get?

OpenStudy (anonymous):

Are you going to kill me?

OpenStudy (mertsj):

Ok. Let's just do the left side What is -6(5-2z) ?

OpenStudy (anonymous):

-30

OpenStudy (mertsj):

-6(5-2z)=-30+12z

OpenStudy (mertsj):

No I'm not going to kill you I'm going to revoke your birthday.

OpenStudy (anonymous):

Revoke my birthday might as well just shoot me?

OpenStudy (anonymous):

Solve. Please show all of your work. If the first and third of three consecutive odd integers are added, the result is 69 less than five times the second integer. Find all three integers

OpenStudy (mertsj):

Do you agree that -6(5-2z)=-30+12z ???

OpenStudy (anonymous):

I been agree with you.

OpenStudy (mertsj):

So are we done with that other problem and you are going to stop posting it?

OpenStudy (anonymous):

Yes I will stop posting it up and we are definitely done with that, Can we proceed?

OpenStudy (anonymous):

Solve. Please show all of your work. If the first and third of three consecutive odd integers are added, the result is 69 less than five times the second integer. Find all three integers

OpenStudy (mertsj):

If the first and third of three consecutive odd integers are added, the result is 69 less than five times the second integer. Find all three integers

OpenStudy (mertsj):

Let the three consecutive odd integers be: x x+2 x+4

OpenStudy (mertsj):

x+x+4=5(x+2)-69

OpenStudy (mertsj):

Do you agree with that equation?

OpenStudy (anonymous):

Yes

OpenStudy (mertsj):

Can you solve it?

OpenStudy (anonymous):

Im slow I can't solve it?

OpenStudy (mertsj):

\[2x+4=5x+10-69\]

OpenStudy (mertsj):

\[2x+4=5x-59\]

OpenStudy (mertsj):